Consider the following information of Jones Corporation over four years Sales Cost of goods sold Gross profit Gross profit percentage 2022 202120202019 $10,000 $9,000 $? $7,000 ? 6,840 6,160 2,500 ? 1,840 22% Required: 1. Calculate the missing amounts for each year. 2. What does this information indicate about the company?
Answer #1
2022 2021 2020 2019
Sales 10000 9000 8000 7000
COGS 7500 6840 6160 5460
Gross Profit 2500 2160 1840 1540
Gross Profit % 25% 24% 23% 22%
Notes:
a)1000-2500=7500
b)2500/10000=25%
c)9000-6840=2160
d)2160/9000=24%
e)6160+184=8000
f)1840/8000=23%
g)7000*22%=1540
h)7000-1540=5460
2
Gross profit shows increasing trend
meaning the company is earning healthy profits and more so
an increasing trend on its profits from operations